# Remote Work and Women’s Labor Supply: The New Gender Division at Home

### v1

@@ 2.2 Sample Restrictions
We restrict our primary analysis, based on the CPS, to dual-parent households in which both the husband and wife reside together and have at least one child of child-rearing age. Specifically, we focus on households with children between the ages of 0 and 8. We exclude households with more than four children. Additionally, we restrict the sample to women aged 20 to 50 and men aged 18 to 64.
In terms of time coverage, we use both the CPS basic monthly and March files spanning 2017 to 2024. The March files contain both weekly labor force variables—comparable to those in the basic monthly files—and additional variables such as annual employment, weeks worked, and wage income, which refer to the previous calendar year. As a result, the effective data coverage for regressions using these annual March files variables is 2016 to 2023.111111Note that occupation in the March files is measured at the time of the survey, while some labor market outcomes—such as annual employment and earnings—refer to the previous year. Therefore, in regressions using these outcomes, we implicitly assume that individuals’ occupation at the time of the survey is the same as in the previous year, which may introduce some measurement error.
To avoid conflating temporary COVID-related labor market disruptions with the long-run shift in remote work, we exclude data from March 2020 to March 2023 in analyses using the CPS basic monthly and March files outcomes referring to the prior year. For regressions using only contemporaneous March CPS files variables—such as weekly labor force status—we exclude data from March 2020 through December 2022, allowing us to include the 2023 March CPS sample and thereby add more post-COVID observations. [^12] The federal COVID-19 Public Health Emergency officially ended on May 11, 2023. For more details, see this link. In the robustness section, we show that our results are not sensitive to alternative definitions of the post-COVID period.
@@ 2.3 Treated and Control Group Definitions
We use job postings data from Hansen et al. (2023) to construct occupation-level measures of the change in WFH probability before and after the COVID-19 pandemic. Specifically, we compute the difference in the share of hybrid and remote-work job postings between the post-pandemic period (2023–2024) and the pre-pandemic period (2019). [^13] As an alternative, we also construct a measure based on the change in WFH share between two specific months: July 2023 and July 2019.
@@
We convert occupation codes from the original SOC 3-digit classification used in the job postings data to the occupation codes used in the CPS. Occupations with a WFH probability change at or above the median are classified as treated, and those below the median as control. [^14] As a robustness check, we alternatively define treatment based on the median split within the original SOC 2018 3-digit classification. These occupation-level measures capture the pandemic-induced shift in WFH feasibility and are merged to our household survey data based on the husband’s occupation at the time of the survey. [^15] Ideally, we would match individuals to their pre-pandemic occupation to avoid concerns about endogenous selection into more WFH-intensive occupations on the basis of unobservables.
@@
Appendix Table B.1 lists the 15 most common occupations in the treated and control groups. As expected, occupations that experienced the largest increases in WFH probability after COVID-19 include those in the tech sector, law, and management,such as computer and software engineers, lawyers, and various types of managers. In contrast, occupations with minimal changes in WFH feasibility are concentrated in sales, construction, and retail, as well as in sectors like healthcare (e.g., physicians and surgeons) and education (e.g., teachers), where in-person presence remains a key job requirement.
@@
Appendix Table B.2 and Table B.3 present descriptive statistics for baseline labor market outcomes and household-level covariates, respectively. On average, treated households are more likely to have both spouses employed, [^16] For spouse employment and labor force participation measured in the previous week, this pattern is mechanical: the variable equals 1 for all observations because occupation—and hence treatment status—is only recorded for individuals who are currently employed or in the labor force. with both husbands and wives working at higher rates than in control households. As expected, treated households are more likely to have both spouses highly educated and less likely to have attained only a high school degree or less. Treated wives are also slightly older, on average, than their counterparts in the control group. In terms of family composition, treated households tend to have fewer children and younger children on average. Additionally, labor market outcomes for women in the treated group exhibit larger average changes post-COVID compared to women in the control group, though this pattern is purely descriptive. In all regressions, we control for the education levels of both spouses, the wife’s age, and the number of children in the household. In robustness checks, we additionally include controls for the age of the youngest and oldest child (see Table B.16).
As a robustness exercise, we additionally define treatment using the teleworkability measure developed by Dingel and Neiman (2020), which classifies occupations based on their feasibility to be performed remotely. Occupations at or above the median in this measure are classified as treated, and those below as control. Unlike our baseline, which relies on realized job postings and thus reflects employer demand, this measure is constructed from occupational characteristics and reflects potential—rather than actual—telework feasibility. [^17] Importantly, this classification does not account for demand-side frictions or social norms that may affect the practical implementation of remote work. For instance, while Dingel and Neiman (2020) classify teachers as capable of WFH, post-pandemic trends suggest that in-person instruction remains the preferred mode, largely due to parental demand (Barrero et al., 2023). Crucially, both this and our baseline measure are based on occupational characteristics and do not rely on actual WFH take-up by husbands, which helps avoid concerns about endogenous sorting or selection based on realized behavior (Mas and Pallais, 2017).
@@ 3 Empirical Strategy
3.1 Baseline Difference-in-Differences
We exploit variation in the change in work-from-home (WFH) probability across occupations induced by the COVID-19 pandemic. Our identification strategy compares labor market outcomes of women whose husbands were employed in occupations that experienced larger shifts in WFH probability, before and after the pandemic. As explained in the previous section, change in WFH exposure is collapsed at the occupation level and reflects the difference in demand-driven WFH probabilities pre- and post-COVID.
@@ 3.1 Baseline Difference-in-Differences
$\begin{split}y_{i(H),s,t}=\;&f_{s,t}+f_{o(H)}\;+\\ &\beta\times 1\left(o(H)\in\{\text{High $\Delta$WFH Occupations}\}\right)\times 1(t>\text{March 2023})\;+\\ &X_{i(H),t}\delta+\epsilon_{i(H),s,t}\end{split}$ (1)
where $y_{i(H),s,t}$ denotes a labor market outcome for woman $i$, married to husband $H$, observed in state $s$ and time tt. [^18] Time is defined at the month-year level in CPS, ATUS, and SIPP. In ACS, where month of interview is unavailable, time is defined at the year level. The interaction term captures the treatment effect of interest: it equals one for observations after March 2023 in which the husband’s occupation $o(H)$ belongs to the set of occupations that experienced an above-median increase in the probability of WFH between the pre- and post-pandemic periods. [^19] In the ACS sample, where data are available only at the annual level from 2017 to 2023, we define the post-period as years 2022 and 2023, due to the absence of month-level timing and in order to have a sufficient sample size to run our regressions.202020 In the ATUS sample, where we have data from 2017 to 2024, and it is a subsample from CPS, we define March 2022 as the starting point of the post-treatment period to have sufficient sample size.
The control vector $X_{i(H),t}$ includes the woman’s age, education, and number of children, as well as the husband’s education. All regressions include state-by-time fixed effects $f_{s,t}$ to account for local labor market shocks and policy variation, and occupation fixed effects $f_{o(H)}$ to absorb time-invariant differences across husband’s occupations.
@@
We also conduct industry-level specifications, where we construct a treatment indicator based on whether the husband’s Census 2012 industry ($n(H)$) experienced an above-median increase in WFH probability. In these regressions, we replace the occupation-level treatment term in Equation (1) with the industry-level indicator interacted with the post-COVID period. Additionally, we estimate a combined occupation–industry specification that includes three interaction terms: (i) the occupation-based treatment indicator interacted with the post-COVID period, (ii) the industry-based treatment indicator interacted with the post-COVID period, and (iii) their interaction, also interacted with the post-COVID dummy. This approach allows us to separately identify the effects of occupational and industry WFH exposure, as well as any potential complementarities between the two. All specifications include state-by-time and spouse occupation and/or industry fixed effects, along with the standard set of demographic controls.
@@
Under the standard parallel trends assumption—namely, that in the absence of the pandemic, the labor market outcomes of women whose husbands worked in occupations that experienced large increases in WFH feasibility would have followed similar trends to those whose husbands worked in occupations with minimal changes—the coefficient $\beta$ identifies the causal effect of interest. Specifically, $\beta$ captures the intra-household spillover effect of the husband’s exogenous shift in WFH probability, induced by the pandemic, on the wife’s labor market outcomes.
We conduct a range of robustness checks to support the validity of our identification strategy. These include tests for selection and group composition, the exclusion of influential occupation groups to ensure that results are not driven by specific job categories, alternative specifications of fixed effects and controls, and several placebo exercises. Across all these tests, our results remain consistent, as discussed in detail in Section 5.
It is important to note that, due to data limitations, we are unable to fully isolate spillover effects from potential direct effects. In most surveys, occupational information is only recorded for respondents who are currently employed. As a result, we cannot systematically observe or control for the wife’s own WFH exposure or estimate the interaction between the husband’s and wife’s occupational flexibility. [^21] See Appendix C for a detailed discussion of the identification challenges posed by the lack of occupational data for non-employed individuals and the implications for isolating intra-household spillover effects from potential direct effects of the wife’s own WFH exposure.
In the next section, we present estimates of the coefficient of interest ($\beta$) from this difference-in-differences specification, focusing on a range of labor market outcomes at both the weekly and annual level.
3.2 Dynamic Difference-in-Differences
@@
$y_{i(H),s,t}=$ $f_{s,t}+f_{o(H)}\;+$ (2)
$\sum_{\begin{subarray}{c}y=2017\\ y\neq 2019\end{subarray}}^{2024}\beta^{y}\cdot 1\left(o(H)\in\{\text{High $\Delta$WFH Occupations}\}\right)\cdot 1\left(\text{year}(t)=y\right)$
$+X_{i(H),t}\delta+\epsilon_{i(H),s,t}$
where 2019 serves as the omitted (base) year. The coefficients $\beta^{y}$ capture the differential evolution of outcomes for women whose husbands were in occupations with large WFH shifts, relative to those in occupations with smaller or no changes, for each year from 2017 to 2024 (excluding 2019). All notation follows as in Equation (1).
We plot the estimated $\beta^{y}$ coefficients to visually assess the plausibility of the parallel trends assumption and to explore the dynamic pattern of the estimated effects.
@@ 4.1 Baseline Difference-in-Differences Estimates
Our main identification strategy exploits how occupation-level shocks in women’s spouses increased women’s labor market participation by plausibly increasing spouse’s time availability for household production. Consequently, our first stage analysis consists of verifying that large shocks on occupation-level remote work options led to sizable increase in uptake of remote work for men.
Because the CPS started asking monthly about remote work on 2023 after the COVID-19 pandemic, we cannot use our pre-period sample to estimate the baseline specification. However, the CPS questionnaire asks respondents whether they used to work from home in their occupation in the months preceding the start of the pandemic. With this recall data, we are able to build a pseudo-panel dataset with a single post-Covid and pre-Covid observation per household-month. Both observations have the same covariates and vary only in their timing status and whether or not they worked from home using the actual post-Covid and recalled pre-covid responses.
In Table 1, we present our first stage results using this pseudo-panel CPS data. We implement our baseline specification from Equation (1) and an alternative panel specification that uses household- and time- fixed effects. Both specifications show a very large increase in actual uptake of remote work of around 17 to 20 percentage points in our sample of male spouses. With a pre-pandemic average WFH rate of 10%, rates have tripled in our sample households. This increase roughly matches the large increases in share of remote work job vacancies for high exposed occupations that we documented in our motivating Figure 1.
Further, we report estimates separately for the subsample of men that are employees and the self-employed. As expected, WFH take-up increased significantly more among employees, who experienced a threefold rise relative to pre-COVID levels, compared to the self-employed, whose WFH rates increased by approximately 50 percent from a higher baseline of 16 percent. It is worth remarking that even for the self-employed the increase is large and highly significant. This motivates our choice to keep these households in our sample for the main analysis.
@@
Table 1: First-Stage Results: Effect of Occupation-Level WFH Shocks on Husbands’ Remote Work Uptake
Cross-section Panel
Spouse Restriction: All Employee Self-employed All Employee Self-employed
$Exposed_{o(i)}\times Post_{t}$ 0.203*** 0.217*** 0.087*** 0.175*** 0.190*** 0.079***
@@
Mean pre-treatment 0.106 0.099 0.161 0.106 0.099 0.161
Obs 126,240 111,135 15,049 124,549 109,577 14,702
$R^{2}$ 0.30 0.32 0.32 0.66 0.67 0.66
  • Notes: We regress men’s self-reported telework status on their treatment status, using both cross-sectional and pseudo-panel specifications. The data comes from our household sample in the 2023–2024 monthly CPS. Respondents are asked whether they currently work from home (post-COVID) and whether they did so in February 2020 (pre-COVID), allowing us to construct a two-period pseudo-panel under the assumption that respondents remained in the same occupation. In the first three columns, we estimate cross-sectional regressions that include month-year, state, and spouse-occupation fixed effects. In the final three columns, we exploit the panel structure of the recalled data by including household fixed effects instead of occupation fixed effects. For each specification, we report results for the full sample, the subset of households where the male spouse is an employee, and the subset where he is self-employed.
Effects on Labor Market Outcomes
Our hypothesis is that the expansion of remote work opportunities for men relaxes constraints on household production, both in terms of time and physical presence, thereby enabling their spouses to increase labor market engagement while raising children. Table 2 presents our main estimates from Equation (1), quantifying the effect of increased WFH exposure for husbands on their wives’ labor market outcomes. The results indicate a sizable and statistically significant increase of approximately 2.5 percentage points in women’s probability of employment and labor force participation, relative to a pre-treatment mean of 67% — equivalent to a 4% increase.
While the effects on usual weekly hours, months worked, and part-time status are less precisely estimated, we find a meaningful decline in the probability of part-time work using the March files, consistent with an increase in labor supply intensity. In addition, we estimate an 8% increase in women’s wages, suggesting potential productivity gains or access to better job matches facilitated by the reallocation of household responsibilities. To explore treatment heterogeneity, Table B.4 reports estimates where we discretize the treatment variable into quartiles of the husband’s occupation-level WFH exposure. The first quartile corresponds to occupations with the smallest changes in WFH probability, while the fourth captures those with the largest. As expected, the effects are concentrated among households where the husband’s occupation falls into the top two quartiles of the WFH exposure distribution.
Table 2: Effect of Husbands’ WFH Shock on Wives’ Labor Market Outcomes
Monthly CPS sample March CPS sample
Employed LFP Usual work Part-time Employed LFP Part-time Months Log wages
hours status Status Worked
$Exposed_{o(H)}\times Post_{t}$ 0.024** 0.027*** 0.217 0.004 0.035** 0.027* -0.030* 0.120 0.078**
(0.011) (0.010) (0.197) (0.008) (0.015) (0.014) (0.016) (0.095) (0.036)
Mean pre-treatment 0.672 0.686 36.631 0.313 0.665 0.708 0.247 10.897 9.992
Obs 252,309 252,309 162,882 161,095 38,637 38,637 25,920 25,920 25,920
$R^{2}$ 0.12 0.12 0.06 0.06 0.12 0.13 0.07 0.07 0.17

### v2

@@ 3.2 Sample Restrictions
We restrict our primary analysis, based on the ACS, to dual-parent households in which both the husband and wife reside together and have at least one child of child-rearing age. Specifically, we focus on households with children between the ages of 0 and 8 and exclude households with more than four children. Additionally, we restrict the sample to women aged 20 to 50 and men aged 18 to 64.141414We adopt standard age restrictions for women, as is common in the literature (Kleven, 2024). We additionally impose age restrictions on men to ensure they are of typical labor-force age.
In terms of time coverage, we use ACS files spanning 2017 to 2023, the most recent year of data. The files contain both weekly labor force variables — comparable to those in the basic monthly files of CPS — and additional variables such as annual employment, weeks worked, usual weekly hours worked, part-time status and earnings, which refer to the previous 12 months. As a result, the effective data coverage for regressions related to these annual outcomes is 2016 to 2022.151515Note that occupation is measured/reported at the time of the survey, while annual outcomes refer to the previous 12 months. Therefore, in regressions using these outcomes, we implicitly assume that individuals’ occupation at the time of the survey is the same as in the previous 12 months, which may introduce some measurement error.
@@ 3.3 Treated and Control Group Definitions
We use job postings data from Hansen et al. (2023) to construct occupation-level measures of the change in WFH probability before and after the COVID-19 pandemic. Specifically, we compute the difference in the share of hybrid and remote-work job postings between the post-pandemic period (2023–2024) and the pre-pandemic period (2019). [^16] As an alternative, we also construct a measure based on the change in WFH share between two specific months: July 2023 and July 2019. Figure A.1 shows the distribution of our WFH exposure measure.
@@
We convert occupation codes from the original SOC 3-digit classification used in the job postings data to the occupation codes used in the ACS. Occupations with a WFH probability change at or above the median are classified as treated, and those below the median as control. [^17] As a robustness check, we alternatively define treatment based on the median split within the original SOC 2018 3-digit classification. These occupation-level measures capture the pandemic-induced shift in WFH feasibility and are merged to our household survey data based on the husband’s occupation at the time of the survey. [^18] Ideally, we would match individuals to their pre-pandemic occupation to avoid concerns about endogenous selection into more WFH-intensive occupations on the basis of unobservables.
@@
Appendix Table B.1 lists the 15 most common occupations in the treated and control groups. As expected, occupations that experienced the largest increases in WFH probability after COVID-19 include those in the tech sector, law, and management,such as computer and software engineers, lawyers, and various types of managers. In contrast, occupations with minimal changes in WFH feasibility are concentrated in sales, construction, and retail, as well as in sectors like healthcare (e.g., physicians and surgeons) and education (e.g., teachers), where in-person presence remains a key job requirement. We use both the 2019 Occupational Employment Statistics (OES) data (Table B.2) and 2019 ACS data (Table B.3) to compute characteristics of treated and control occupations. As expected, treated occupations are associated with higher wages, both on average and at any percentile of the distribution, and tend to employ more educated workers with longer schooling and longer work hours. Somewhat surprisingly, women appear slightly more represented in treated occupations in the ACS sample, although the difference is modest (about 3 percentage points) and not economically significant.
Using the flexibility indexes from Goldin (2014) - updated with the latest O*NET data - we find little correlation between our treatment definition and these flexibility measures (Figure A.2, Figure A.3), easing concerns about pre-existing occupational differences. As shown in Table B.4, treated occupations are somewhat more WFH-suitable (e.g., less physical proximity, more irregular schedules) but not necessarily more “flexible” in the sense of Goldin’s indexes, which are statistically higher for treated occupations. [^19] Higher index values indicate lower worker substitutability and greater job flexibility. This distinction underscores that while remote work is related to flexibility, the two are not equivalent.
@@
Appendix Table B.5 and Table B.6 present descriptive statistics for baseline labor market outcomes and household-level covariates, respectively. On average, treated households are more likely to have both spouses employed, [^20] For spouse employment and labor force participation measured in the previous week, this pattern is mechanical: the variable equals 1 for all observations because occupation—and hence treatment status—is only recorded for individuals who are currently employed or in the labor force. with both husbands and wives working at higher rates than in control households. As expected, treated households are more likely to have both spouses highly educated and less likely to have attained only a high school degree or less. Treated wives are also slightly older, on average, than their counterparts in the control group. In terms of family composition, treated households tend to have fewer children and younger children on average. Additionally, labor market outcomes for women in the treated group exhibit larger average changes post-COVID compared to women in the control group, though this pattern is purely descriptive. In all regressions, we control for the education levels of both spouses, the wife’s age, and the number of children in the household.
@@ 4 Empirical Strategy
Baseline Difference-in-Differences
We exploit variation in the change in WFH probability across occupations induced by the COVID-19 pandemic. Our identification strategy compares labor market outcomes of women whose husbands were employed in occupations that experienced larger shifts in WFH probability, before and after the pandemic. As explained in the previous section, change in WFH exposure is collapsed at the occupation level and reflects the difference in demand-driven WFH probabilities pre- and post-COVID.
@@
$y_{ist}=(\beta_{1}Post_{t}+\beta_{2}Post_{t}\times During_{t})\times\text{High-}\Delta\text{WFH}_{\,o(h(i))}+X^{\prime}_{it}\delta+\gamma_{st}+\Theta_{o(h(i))}+\varepsilon_{ist}.$ (4)
where $y_{ist}$ denotes a labor market outcome for woman $i$, observed in state $s$ and time tt. [^21] Time is defined at the month-year level in CPS, ATUS, and SIPP. In ACS, where month of interview is unavailable, time is defined at the year level. $\text{High-}\Delta\text{WFH}_{\,o(h(i))}$ is a dummy for treated occupations, whereas $During_{t}$ and $Post_{t}$ are time-dummies for the years 2020-2021 and for years 2020 onwards, respectively. The coefficient on the first interaction term, $\beta_{1}$, captures the treatment effect of interest: it equals one for observations after March 2020 in which the husband’s occupation $o(h(i))$ belongs to the set of occupations that experienced an above-median increase in the probability of WFH between the pre- and post-pandemic periods. The coefficient on the second interaction termm, $\beta_{2}$, captures the effect only during the covid period, from 2020 to 2021.  [^22] In the ACS sample, where data are available only at the annual level from 2017 to 2023, we define the post-period as years from 2020 to 2023, due to the absence of month-level timing and in order to have a sufficient sample size to run our regressions. The during period is set for the 2020 and 2021 period. In the ATUS sample, a subsample from CPS, where we have data from 2017 to 2024, we define March 2020 as the starting point of the post-treatment period and March 2022 as the end date of the during period, in order to have sufficient sample size.
The control vector $X_{i,t}$ includes the woman’s age, education, and number of children, as well as the husband’s education. All regressions include state-by-time fixed effects $\gamma_{s,t}$ to account for local labor market shocks and policy variation, and occupation-of-husband fixed effects $\Theta_{o(h(i))}$ to absorb time-invariant differences across husband’s occupations.
@@
We also conduct industry-level specifications, where we construct a treatment indicator based on whether the husband’s Census 2012 industry ($n(h(i))$) experienced an above-median increase in WFH probability. In these regressions, we replace the occupation-level treatment term in Equation (4) with the industry-level indicator interacted with the post and during-COVID periods. As with the main specification, we include state-by-time and spouse occupation and/or industry fixed effects, along with the standard set of demographic controls.
@@
Under a standard parallel trends assumption — namely, that in the absence of the pandemic, the labor market outcomes of women whose husbands worked in occupations that experienced large increases in WFH probability would have followed similar trends to those whose husbands worked in occupations with minimal changes — the coefficient $\beta_{1}$ identifies the causal effect of interest. Specifically, $\beta_{1}$ captures any effect of the husband’s exogenous shift in WFH probability, induced by the pandemic, on the wife’s labor market outcomes.
Disentangling Spillovers from Matching Effects
We categorize the effects on women’s labor outcomes into two categories. First, spillover effects operating through the husband’s increased WFH exposure — for instance, via reallocation of childcare or household production — which can shift the wife’s labor supply even holding her own WFH exposure fixed. Second, correlated direct effects arising from assortative matching in marriage markets, whereby wives of men in highly WFH-feasible occupations may themselves hold (or seek) WFH-feasible jobs, mechanically correlating the husband’s exposure with the wife’s own exposure. A key empirical limitation is that most household surveys record occupation only for the employed; the wife’s WFH exposure is therefore missing precisely when she is non-employed. This constrains our ability to condition on, or interact with, the wife’s occupation in the full sample. [^23] See Appendix C for a detailed discussion of these measurement and identification issues and their implications for separating spillovers from direct effects.
Given these constraints, in Section 5 we adopt complementary strategies to separate and assess the contributions of spillover effects (operating through the husband’s WFH exposure) and correlated direct effects (arising from assortative matching) to the overall estimated effect.
Dynamic Difference-in-Differences
@@
$y_{ist}=\sum_{t\neq 2019}\beta_{t}\,\text{High-}\Delta\text{WFH}_{\,o(h(i))}\times D_{t}+X^{\prime}_{it}\delta+\gamma_{st}+\Theta_{o(h(i))}+\varepsilon_{ist}.$ (5)
where 2019 serves as the omitted (base) year. The coefficients $\beta_{t}$ capture the differential evolution of outcomes for women whose husbands were in occupations with large WFH shifts, relative to those in occupations with smaller or no changes, for each year from 2016 to 2023 (excluding 2019). All notation follows as in Equation (4).
@@ 5 Main Results
Our main identification strategy exploits how occupation-level shocks in women’s spouses increased women’s labor market participation by plausibly increasing spouse’s time availability for household production. Consequently, our first stage analysis consists of verifying that large shocks on occupation-level remote work options available to fathers of young children in our ACS sample led to sizable increase in their uptake of remote work. By and large we can confirm this claim. Figure 2 shows that men in treated occupation increase their uptake of remote work by 20 p.p with respect to the control group and relative to pre-COVID year 2019. Although the effect decreases over time, it clearly does not to return to pre-covid level and is still 10 p.p.in 2023242424We use as measure of remote work in ACS the following question: How did you usually get to work LAST WEEK. Mark (X) ONE box for the method of transportation used for most of the distance. As discussed in Buckman et al. (2025), this question design yields a measure for the incidence of fully remote work, and aligns with comparable measures from the Survey on Working Arrangements SWAA and other surveys.. With a pre-pandemic average WFH rate of 4%, rates for men have tripled well after the pandemic in our sample households. This increase roughly matches the large increases in share of remote work job vacancies for high exposed occupations that we documented in Figure 1.
Figure 2: Dynamic Effect of Husbands’ WFH Shock on Own Remote Work Uptake
Notes: The figure plots coefficients from a dynamic difference-in-differences specification comparing husbands who were in occupations with large WFH shocks to those in occupations with smaller changes. The omitted year is 2019. All regressions include state-by-time and husband-occupation fixed effects, and control for wife’s age, number of children, and education levels of both spouses. Confidence intervals are shown at the 95% level and are based on standard errors clustered at the husband-occupation level.
Table 1 reports spouses’ WFH uptake estimates from the baseline DiD in (4). On average, husbands’ WFH probability increases by about 13 percentage points (pp) in the post-pandemic period, with larger gains concentrated in the COVID years (2020–2021), as shown by the post-Covid and treatment interaction. We also break down WFH uptake by employment status of the husband. The increase is markedly larger for employees — roughly a doubling relative to pre-COVID levels — whereas the self-employed show a smaller but still sizeable rise of about 40 percent from a higher pre-period baseline of 14 percent. In the final column, we examine commuting time (minutes): men in more exposed occupations reduce commuting by about 5 and a half minutes, roughly 18% of the pre-treatment mean.
Overall, the increase in WFH uptake among men was as large as the shock in available jobs with remote work options available in their occupations. Although the effects are much smaller by 2023 - which corresponds with more companies increasingly mandating return to office among their employees - the shocks have had lasting effects, well after Covid. [^25] ACS captures only fully remote work, but nowadays hybrid options are more prevalent. Moreover, these shocks did not immediately increase men’s intensive margin of labor supply. In Table B.7, we find that the remote work shock did not increase working hours or wages among men, and if anything, men spend marginally less hours of the week working. From the household’s perspective, the shock initially led to a higher and long-term availability of men at households with young children.
@@
Table 1: Effect of Occupation-Level WFH Shocks on Husbands’ Remote Work Uptake
Fully Remote Work Commuting
All Employee Self-employed Time To Work
$Exposed_{o(h(i))}\times Post_{t}$ 0.133*** 0.139*** 0.053*** -5.384***
(0.017) (0.018) (0.012) (0.647)
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ 0.038*** 0.039*** 0.017 -1.247***
(0.007) (0.007) (0.019) (0.265)
Mean pre-treatment 0.044 0.038 0.136 29.688
$\beta_{1}+\beta_{2}=0$ (p-val) 0.000 0.000 0.000 0.000
Obs 704,497 659,974 44,491 704,408
$R^{2}$ 0.17 0.18 0.14 0.06
  • Notes: Estimates from Eq. (4) using husbands’ self-reported telework status as the outcome. The sample is the ACS, 2017–2023. Outcomes refer to the week prior to the interview. We report results for (i) the full sample, (ii) households in which the husband is an employee, and (iii) households in which he is self-employed. The WFH measure is constructed from the commuting-mode question—“How did you usually get to work LAST WEEK? Mark (X) ONE box for the method of transportation used for most of the distance.”—and is interpreted as the incidence of fully remote work; see Buckman et al. (2025). All specifications control for the wife’s age, education, and number of children, as well as the husband’s education. State-by-time and husband-occupation fixed effects are included. Standard errors are clustered at the occupation level.
Effects on Labor Market Outcomes
We hypothesize that expanded WFH opportunities for men relax constraints on household production — both temporal and physical — thereby easing wives’ participation in the labor market while raising children. Table 2 presents estimates from Eq. (4) and shows that shocks on available remote work options for women’s husbands increase women’s labour supply both in the extensive and intensive margins. [^26] We find similar results if we split occupations by quartiles of the treatment exposure. Effects are concentrated and largely similar among the third and fourth quartiles as shown in Table B.8. We find that the probability a wife is employed increases by about 2.5 percentage points — roughly a 4% rise relative to the 69% pre-treatment mean. Total annual earnings increase by around 5%. Weekly hours rise by approximately 0.5 hours, and the number of weeks worked increases by 0.66. Consistent with these intensive-margin responses, the likelihood of part-time work falls by about 2.5 percentage points (9% relative to the pre-treatment mean), while we find no statistically significant effect on hourly wages. [^27] We also evaluate how our effects would change if we defined treatment at the male spouse’s industry occupation-level. Figure A.4 shows time series evolution of industry-level WFH trends between 2017 and 2023, showing largely analogous effects to the occupation-level trends shown in Figure 1. In Table B.10, we present estimates of our baseline specification but using the husband’s industry-level WFH exposure as the treatment variable. The results go in the same direction as those based on occupational exposure, but are lower in magnitude and not as significant. This is consistent with our prior that employer industry may not be as precise as occupation to identify available remote work options. Within a given industry or firm - e.g. hospitals or education -, certain jobs are much more likely to be teleworkable than others - e.g. office admin vs health practitioner.
Because log earnings are observed only for individuals with positive earnings, our earnings, hours, and wage regressions are estimated on the sample of individuals observed working in a given year. In a repeated cross section, this sample includes both individuals who were already employed and individuals who enter employment in the post period, though we cannot distinguish between these groups. To relate these intensive-margin estimates to an unconditional earnings concept, we use a standard two-part decomposition. Unconditional mean earnings can be written as $\mathbb{E}[A]=\Pr(E=1)\cdot\mathbb{E}[A\mid E=1]$. Using a first-order approximation, the treatment effect on log mean earnings is
$\Delta\log\mathbb{E}[A]\approx\Delta\log\Pr(E=1)+\Delta\log\mathbb{E}[A\mid E=1].$ (6)
We approximate the first term by $\beta_{E}/\bar{E}$, where $\beta_{E}$ is the estimated employment effect and $\bar{E}$ is the baseline employment rate in the control group, and the second term using the coefficients from the log earnings decomposition estimated among individuals observed working. [^28] Because earnings among working individuals can be written as $A=w\times H\times WW$, where $w$ denotes the hourly wage, $H$ usual hours worked per week, and $WW$ weeks worked during the year, we decompose changes in log earnings conditional on employment as $\Delta\log\mathbb{E}[A\mid E=1]\approx\beta_{\log w}+\beta_{\log H}+\beta_{\log WW}$. When hours or weeks are measured in levels, coefficients are converted to log changes using a first-order approximation evaluated at baseline means.
Applying this decomposition, we find that intensive-margin responses account for approximately 2.9 log points (46%) of the overall earnings effect, reflecting contributions from hourly wages (0.4 log points), usual hours worked (1.3 log points), and weeks worked (1.2 log points), while the extensive margin contributes an additional 3.4 log points (54%) through increased employment. Together, these components imply an overall earnings effect of about 6.3 log points, which exceeds the coefficient from the conditional log earnings regression. This difference is consistent with newly employed individuals having lower earnings than incumbent workers, so that changes in the composition of the employed sample attenuate the estimated effect on mean log earnings. [^29] Because the log earnings regression is conditional on positive earnings, the composition of the employed sample changes when employment increases. In a repeated cross section, the post-period sample of earners includes both pre-existing workers and new entrants, whom we cannot separately identify. As a result, the decomposition is descriptive and need not sum exactly to the coefficient from the conditional log-earnings regression, even though the earnings identity holds at the individual level.
Appendix Table B.9 reports a complementary specification in which earnings and its components are defined for all individuals, with zeros imputed for non-employed individuals. In this case, the earnings identity holds mechanically and the decomposition is exact. The resulting estimates show that the “wage” component closely mirrors the overall earnings effect, reflecting transitions from zero to positive wages associated with increased employment rather than changes in hourly pay among workers, while changes in hours and weeks play a more limited role.
Table 2: Effect of Husbands’ WFH Shock on Wives’ Labor Market Outcomes
Employment Log Earnings Log Wage Usual Weekly Hours Weeks Worked Part-time
Worked Status
$Exposed_{o(h(i))}\times Post_{t}$ 0.024*** 0.052*** 0.004 0.472*** 0.666*** -0.025***
@@
$Exposed_{o(h(i))}\times Post_{t}\times During_{t}$ -0.009* -0.009 -0.005 0.082 -0.236* 0.001
(0.005) (0.012) (0.009) (0.113) (0.120) (0.005)
Mean pre-treatment 0.698 9.956 2.695 36.390 46.060 0.272
$\beta_{1}+\beta_{2}=0$ (p-val) 0.000 0.000 0.857 0.000 0.000 0.000
Obs 729,120 514,151 514,151 514,151 514,151 514,151
$R^{2}$ 0.08 0.15 0.20 0.03 0.03 0.04